HCO^{+} IN THE ENVELOPES OF OXYGEN-RICH STARS: EVALUATING THE ROLE OF ION-MOLECULE CHEMISTRY IN CIRCUMSTELLAR GAS

\underline{R.L. PULLIAM}, L. M. ZIURYS, University of Arizona, Steward Observatory, 933 N. Cherry Ave, Arizona.

The role of ion-molecule chemistry is not well understood. Theory has predicted HCO^{+} as one of the most abundant detectable ions in oxygen-rich envelopes. However, there has never been a conclusive detection of HCO^{+} in any circumstellar sources. Here we present the results of millimeter and submillimeter observations of the cation HCO^{+} around evolved oxygen-rich stars. In recent observations, HCO^{+} has been detected in the circumstellar shell of the oxygen-rich supergiant VY Canis Majoris. Further observations were conducted around the oxygen-rich stars IK Tau, NML Cyg, W Hya and TX Cam. Abundances of this molecule in these sources and implications for circumstellar chemistry will be discussed.